# Model Predictive and Reinforcement Learning Methods for Active Flow Control of an Airfoil with Dual-point Excitation of Plasma Actuators

### v1

@@ Abstract
This paper presents an analysis of Model Predictive Control (MPC) and Reinforcement Learning (RL) approaches for active flow control over a NACA (National Advisory Committee for Aeronautics) 4412 airfoil around the static stall condition at a Reynolds number of ${4\times 10^{5}}$. The Reynolds Averaged Navier-Stokes (RANS) equations with the Scale-Adaptive Simulation (SAS) turbulence model are utilized for the numerical simulations. The dielectric barrier discharge (DBD) plasma actuators were employed in dual excitation mode for flow separation control. The study systematically evaluates adaptive MPC, temporal difference reinforcement learning (TDRL), and deep Q-learning (DQL) based on optimizing the excitation frequency and expediting the time to identify stable conditions. Moreover, an integrated approach that combines signal processing with DQL is examined. The results demonstrate that while MPC and RL significantly improve flow control, RL approaches offer superior adaptability and performance. In optimal conditions, a lift coefficient of around 1.619 was achieved within less than 2.5 seconds with an excitation frequency of 100 or 200 Hz. This research highlights that RL-based approaches could perform better in flow control applications than MPC.
@@
journal: arXiv
[…]
[inst1]organization=Faculty of Aerospace Engineering, Delft University of Technology,addressline=Kluyverweg 1, city=Delft, postcode=2629 HS, country=Netherlands
[…]
[inst2]organization=Department of Aerospace Engineering, Sharif University of Technology,addressline=Azadi Ave., city=Tehran, postcode=1458889694, country=Iran
[…]
[inst3]organization=College of Interdisciplinary Science and Technology, University of Tehran,addressline=Northern Kargar St., city=Tehran, postcode=1439957131, country=Iran
@@ 1 Introduction
Closed-loop flow control (CFC) methods have emerged as a promising solution for mitigating the adverse hydrodynamic and aerodynamic effects. These methods dynamically monitor and influence the flow field, based on feedback from real-time flow conditions by utilizing sensors and actuators [1, 2]. Implementing CFC methods provides more efficient and adaptable hydro/aerodynamic systems in various engineering fields. In aviation, CFC methods can enhance aircraft maneuverability, delay stall, and improve fuel efficiency [3]. The automotive industry also benefits from these techniques by reducing drag and improving vehicle stability [4]. A critical area where CFC methods demonstrate their value is in managing flow separation, a non-linear phenomenon characterized by the detachment of the boundary layer from the surface, leading to increased drag and reduced lift and aerodynamic efficiency. Thus, research in closed-loop flow separation control has increasingly centered on novel approaches to improve aerodynamic performance. In the following, some of the most relevant works are reviewed.
@@
The CFC methods can be categorized into model-based and data-driven approaches. Model-based methods have greatly enhanced the development of flow control strategies. Proctor et al. [5] expanded dynamic mode decomposition (DMD) to incorporate control effects, resulting in precise input-output models for complex, high-dimensional systems, referred to as dynamic mode decomposition with control (DMDc). Addressing uncertainties in system identification, Dovetta et al. [6] presented a hybrid framework combining perturbative methods and Monte Carlo techniques, efficiently estimating the statistical properties of identified systems. Huang and Kim [7] tackled separated flows using a linear optimal control approach based on system identification, successfully reducing separation bubble size. Lastly, Obeid et al. [8] used the NARMAX identification technique and Proportional–Integral (PI) controller for closed-loop control of flow separation over a NACA 0015 airfoil, significantly improving the lift coefficient and showcasing advanced system identification methods.
@@
Data-driven approaches and deep reinforcement learning (DRL) have recently attracted much attention in the study of flow separation control [9, 10]. Vignon et al. [11] provided an overview of DRL frameworks in active flow control, identifying limitations and future directions. Shimomura et al. [12] showed the effectiveness of a CFC system using a deep Q Network (DQN) on a NACA 0015 airfoil, achieving higher control gains accompanied by the attached flow. A RL-based method for actuator selection was introduced by Paris et al. [13] to optimize control performance of laminar flow separation around an airfoil. Additionally, Vinuesa et al. [14] and Varela et al. [15] reviewed DRL applications in flow control, with the aim of highlighting the potential for drag reduction of wings and circular cylinders.
DRL has been successfully applied to Active Flow Control (AFC) in turbulent conditions, achieving a 30% drag reduction at Reynolds (Re) number of 1000 by reducing turbulent fluctuations and elongating the recirculation bubble [16]. In varying Re number flows around a square cylinder, DRL’s adaptability was demonstrated with drag reductions up to 47%, effectively suppressing vortex shedding across different Re values [17]. In the control of turbulent separation bubbles (TSB), DRL outperformed classical periodic forcing, achieving a 25.3% reduction in TSB length at friction Reynolds number of 180 [18]. At friction Reynolds number of 750, DRL continued to show superiority with an 8.9% reduction in TSB length, while offering smoother control strategies [19].
MPC and RL each have their distinct advantages and challenges. MPC excels at handling constraints effectively by utilizing online optimization, but it can be computationally intensive and challenging to implement. On the other hand, RL methods are well-suited for dynamic, nonlinear environments as they do not require an exact system model, can adapt through experience, and can be implemented online. However, RL typically demands large amounts of data, is sensitive to the design of the reward function and exploration strategy, and involves intricate tuning processes.
@@
Despite considerable progress in flow control methods, the efficiency of CFC methods has not been well documented. Particularly, a significant gap persists in the comparative evaluation of model predictive and RL-based strategies for improving active flow separation control over airfoils using plasma actuators. The existing literature primarily focuses on individual control methods, reinforcement learning algorithms, or specific applications of plasma actuators. However, a comprehensive comparison of these advanced methods to determine their relative effectiveness and potential integration has not been extensively explored. The novelty of this study lies in its systematic comparison of model predictive control and RL-based approaches, including TDRL, DQL, and DQL integrated with signal processing for active flow separation control. This research seeks to offer a more comprehensive understanding of the strengths and weaknesses of each method to advance more effective and efficient flow control strategies.
@@
The rest of the paper is structured as follows: the flow control actuator section introduces the concept of dual-point excitation using plasma actuators. The underlying computational framework, including the governing equations, plasma actuator model, computational grid, and numerical methods used, are explained in the computational procedure section. In the flow control algorithms section, the various algorithms are investigated: adaptive MPC, TDRL, DQL, and signal processing integrated with DQL, each described with their respective algorithms, governing equations, and code implementations. Finally, the Results and Discussion section presents a thorough analysis of the findings, highlighting the comparative performance of the different control strategies and discussing their implications for future research and practical applications.
@@ 2 Flow Control Actuator
Among different AFC devices, DBD plasma actuators have garnered significant attention in flow control applications. Their popularity stems from several key advantages, including being fully electric, the absence of moving parts, rapid response time, low weight, and relatively low power consumption [20, 21]. In the previous work of authors [22], the control of separated shear layers using DBD plasma actuators on a NACA 0015 airfoil was introduced. The findings highlighted the technique’s success in stall prevention and its positive impact on lift-to-drag ratios. Li et al. [23] employed plasma co-flow jets on a NACA 0025 airfoil, which significantly delayed flow separation through enhanced mixing shear layers. Samimy et al. [24] extended active flow control applications to high-speed flows using advancements in plasma actuators to manipulate Kelvin-Helmholtz instabilities. Additionally, Gross and Fasel [25] studied separation control on a NACA 643-618 airfoil using plasma pulsed vortex generator jets, demonstrating high efficiency through linear stability analysis.
@@
Controlling flow separation through plasma actuators, particularly by exciting the separated shear layer, is highly sensitive to the choice of excitation frequency. Research indicates that a free shear layer is most responsive to perturbations near its natural vortex shedding frequency, leading many studies to select frequencies within this range for optimal separation control [27, 28, 29]. However, some research favors higher excitation frequencies [12, 30]. Another critical factor is the location of actuation, typically applied near the leading edge or the suction-side separation point to influence the upper separated shear layer [31]. Ebrahimi and Hajipour [32] introduced the concept of dual-point excitation, as shown in Fig. 1b, where both the suction side and trailing-edge shear layers are simultaneously excited using DBD plasma actuators. This method harnesses the interaction between the two shear layers, offering superior control during static stall conditions by reducing the separation zone and minimizing aerodynamic load fluctuations.
Figure 1: (a) Schematic of a DBD plasma actuator and (b) Schematic of dual-point excitation technique (actuators are indicated with red marks) [31].
@@ 3.1 Governing equations
The numerical flow simulations are performed by analyzing an incompressible viscous Newtonian fluid with constant properties. The governing equations employed for the Scale-Adaptive Simulation (SAS) turbulence model are based on the Reynolds-Averaged Navier-Stokes (RANS) equations, modified to include a scale-adaptive term that allows the model to transition between RANS and Large Eddy Simulation (LES) behavior based on the local flow conditions [33]. The continuity and momentum equations in Cartesian coordinates are:
@@ 3.2 Plasma actuator model
$E_{0}$ represents the maximum electric field strength within the triangular region, and the constants $k_{1}$ and $k_{2}$ are determined to ensure that the field strength at the boundary between plasma and air (along the hypotenuse of the triangle) equals the breakdown value. Consequently, the body force within the plasma zone is given by:
@@
Figure 2: Triangular zone associated with plasma formation [32].
In this research, the parameters of the plasma actuator are consistent with those used by Shyy et al. [34]. The exposed and enclosed electrode lengths are 0.5 mm and 3 mm, respectively. The height of the electrodes is 0.1 mm, and the gap between them in the streamwise direction is 0.25 mm, as illustrated schematically in Fig. 2. The applied voltage is 9 kV (root mean square), with a frequency of 3 kHz. The numerical model of the DBD plasma actuator has been validated by comparing it with Shy et al. [34] work. They assumed that plasma actuators operate in a steady, continuous mode; thus, the body force is modulated based on the duty cycle and excitation frequency to simulate unsteady actions. In this research, a duty cycle of 50% and excitation frequencies ranging from $f=0$ to 400 Hz, corresponding to a normalized frequency of $F^{+}=\frac{f\cdot c}{U_{\infty}}$ in the range of 0 to 11, are employed.
@@ 3.3 Computational grid
The computational test case is based on a NACA 4412 airfoil with a chord length $c=0.4\,\text{m}$, an angle of attack $\alpha=15^{\circ}$, and a Reynolds number $Re=4.0\times 10^{5}$. As depicted in Fig. 3, the computational domain extends $20c$ downstream, $12.5c$ upstream and $12.5c$ in the direction normal to the chord. The airfoil surface is modeled as an adiabatic wall with a no-slip boundary condition. A freestream with a constant velocity of 14.6 m/s enters the domain from the upstream and bottom boundaries, while the downstream and top boundaries are defined as pressure outlets with ambient pressure.
Figure 3: Side view of the computational domain.
@@
Figure 4: Detailed view of computational grid around airfoil and the plasma zones.
Following the previous work of authors [32], the novel concept of dual excitation technique is utilized to control the separated shear layers. Two plasma zones are implemented in the grid: one on the suction side of the airfoil just downstream of the separated shear layer, and the other on the pressure side at the airfoil’s trailing edge. The dimensions of these zones are selected based on the work by Shy et al. [34], with the plasma body force applied only within these regions. The suction side actuator is located at $x=0.35c$, while the pressure side actuator is positioned at $x=0.99c$. The grid’s streamwise resolution in the plasma zones is chosen carefully to satisfy the accuracy requirements for predicting plasma wall jets and flow separation. Fig. 4 illustrates a detailed view of the plasma zones encircled in blue and computational grid around airfoil.

### v2

@@ Abstract
Purpose – This study aims to investigate the effectiveness of Model Predictive Control (MPC) and Reinforcement Learning (RL) approaches for active flow control over a NACA 4412 airfoil near the static stall condition at a Reynolds number of $4\times 10^{5}$. By systematically evaluating these control strategies, the research seeks to address a critical gap in optimizing excitation frequency and improving response time in flow control applications. The study contributes to a deeper understanding of the adaptability and performance of RL-based methods compared to traditional MPC in aerodynamic flow separation control.
Methodology – The study employs a quantitative approach through numerical simulations of the Reynolds Averaged Navier-Stokes (RANS) equations with the Scale-Adaptive Simulation (SAS) turbulence model. Dielectric Barrier Discharge (DBD) plasma actuators, operating in dual-point excitation mode, are utilized for flow separation control. The research evaluates adaptive MPC, temporal difference reinforcement learning (TDRL), and deep Q-learning (DQL) in optimizing excitation frequency and expediting the stabilization process. Additionally, an integrated approach combining signal processing with DQL is examined to enhance control performance.
Findings – This study explores advanced control strategies for optimizing aerodynamic performance by managing flow separation using plasma actuators. We evaluate adaptive MPC, TDRL, DQL, and DQL with signal processing, utilizing dual-point excitation via DBD plasma actuators. Adaptive MPC successfully achieved a target lift coefficient ($C_{l}$) of 1.60 using an excitation frequency of approximately 110 Hz, but struggled to reach higher target $C_{l}$ values near the physical limits. RL methods effectively optimized excitation frequencies, achieving a lift coefficient of approximately 1.62 in under 2.5 seconds with an excitation frequency of 100 or 200 Hz.
Originality – This study presents a novel comparison of RL and MPC methods for active flow control, utilizing DBD plasma actuators to mitigate flow separation and enhance aerodynamic performance. Prior approaches have primarily focused on either MPC or RL independently, often relying on offline learning with separate training and testing phases. In contrast, our research employs an online learning framework, where RL-based techniques such as TDRL, DQL, and signal processing-enhanced DQL dynamically adapt to real-time aerodynamic conditions. By simultaneously evaluating adaptive MPC and RL methods in an online learning setup, we provide new insights into their comparative performance in dynamic environments.
@@
journal: Int. J. Numer. Methods Heat Fluid Flow
Author Accepted Manuscript (AAM) Citation: This is the author accepted manuscript for the article published as: AmirHossein Ghaemi, Abbas Ebrahimi, Majid Hajipour, Seyyed Mohammad Mahdy Shobeiry, Arash Fath Lipaei, “Model Predictive and Reinforcement Learning Methods for Active Flow Control of an Airfoil with Dual-point Excitation of Plasma Actuators”, International Journal of Numerical Methods for Heat & Fluid Flow (Emerald), 2025. DOI: 10.1108/HFF-02-2025-0118 Accepted: 26 June 2025 Version: Author Accepted Manuscript (AAM) – not the final published version. Publisher’s version: The final Version of Record is available at Emerald Insight: link. Abstract page: Emerald Abstract Link. Licence for deposit: This AAM is made available under the Creative Commons Attribution–NonCommercial 4.0 International Licence ($C$ BY-NC 4.0): https://creativecommons.org/licenses/by-nc/4.0/.
@@ 1 Introduction
Closed-loop flow control (CFC) methods have emerged as a promising solution for mitigating adverse hydrodynamic and aerodynamic effects. These methods dynamically monitor and influence the flow field, based on feedback from real-time flow conditions, utilizing sensors and actuators [1, 2]. Implementing CFC methods enables more efficient and adaptable hydro/aerodynamic systems across various engineering fields. In aviation, CFC can enhance aircraft maneuverability, delay stall, and improve fuel efficiency [3]. The automotive industry also benefits from these techniques through drag reduction and improved vehicle stability [4]. In wind energy applications, CFC contributes to increased turbine power output [5], mitigates wake effects to boost overall wind farm performance [6], and helps reduce noise emissions from turbines [7]. Additionally, CFC can alleviate structural loads [8], thereby reducing fatigue and extending the operational lifespan of wind turbines.
A critical area where CFC demonstrates significant value is in managing flow separation, a fundamentally nonlinear and high-dimensional phenomenon that leads to substantial performance losses, such as increased drag, reduced lift, and flow-induced vibrations [9, 10]. Flow separation primarily arises due to adverse pressure gradients that decelerate boundary layer flow, causing detachment and resulting in complex interactions among turbulence, boundary layer dynamics, and unsteady vortical structures [11, 12]. Traditional open-loop control strategies often fall short in addressing these complexities due to their inability to adapt to time-varying flow conditions and disturbances [10]. In contrast, CFC leverages real-time sensor feedback, coupled with actuators and adaptive control algorithms, to dynamically influence flow characteristics, significantly mitigating or even eliminating separated flows [13, 14]. As a result, CFC offers robust and efficient solutions that are essential for managing separated flows in engineering applications. Accordingly, recent research in CFC has increasingly focused on developing novel approaches to improve aerodynamic performance. The following section reviews some of the most relevant contributions in this domain.
@@
Closed-loop flow control methods can be categorized into two main groups: model-based and data-driven approaches. To date, various model-based methods have significantly advanced the development of flow control strategies [15, 16]. Rajesh et al. [17] investigated surface modifications on a circular cylinder with ridges of varying configurations, employing dynamic mode decomposition (DMD) for flow field analysis and a gated recurrent unit (GRU) deep learning model to predict surface pressure distributions efficiently. Proctor et al. [18] extended DMD by incorporating control effects, resulting in precise input-output models for complex, high-dimensional systems, referred to as dynamic mode decomposition with control (DMDc). Addressing uncertainties in system identification, Dovetta et al. [19] proposed a hybrid framework that combines perturbative methods with Monte Carlo techniques to estimate the statistical properties of identified systems efficiently. Huang and Kim [20] tackled separated flows using a linear optimal control strategy based on system identification, achieving a significant reduction in separation bubble size. Moreover, Obeid et al. [21] applied the NARMAX identification technique and a Proportional–Integral (PI) controller for closed-loop control of flow separation over a NACA 0015 airfoil, markedly improving the lift coefficient and demonstrating the effectiveness of advanced system identification methods.
@@
Data-driven approaches and deep reinforcement learning (DRL) have recently attracted much attention in the study of flow separation control [22, 23]. Vignon et al. [24] provided an overview of DRL frameworks in active flow control, identifying limitations and future directions. Shimomura et al. [25] showed the effectiveness of a CFC system using a deep Q Network (DQN) on a NACA 0015 airfoil, achieving higher control gains accompanied by the attached flow. A RL-based method for actuator selection was introduced by Paris et al. [26] to optimize control performance of laminar flow separation around an airfoil. Additionally, Vinuesa et al. [27] and Varela et al. [28] reviewed DRL applications in flow control, with the aim of highlighting the potential for drag reduction of wings and circular cylinders. DRL has been successfully applied to active flow control (AFC) in turbulent conditions, achieving a 30% drag reduction at a Reynolds (Re) number of 1000 by reducing turbulent fluctuations and elongating the recirculation bubble [29]. In varying Re number flows around a square cylinder, DRL’s adaptability was demonstrated with drag reductions up to 47%, effectively suppressing vortex shedding across different Re values [30]. In the control of turbulent separation bubbles (TSB), DRL outperformed classical periodic forcing, achieving a 25.3% reduction in TSB length at a friction Reynolds number of 180 [31]. At a friction Reynolds number of 750, DRL continued to show superiority with an 8.9% reduction in TSB length, while offering smoother control strategies [32].
@@
Despite considerable progress in flow control methods, the efficiency of CFC techniques has not been well documented. The existing literature primarily focuses on individual control methods, reinforcement learning algorithms, or specific applications of flow control actuators. However, a comprehensive comparison of these advanced methods, aimed at determining their relative effectiveness and potential for integration, has not been extensively explored. This study focuses on a systematic comparison of model predictive control (MPC) and RL-based approaches, including TDRL, DQL, and DQL integrated with signal processing, for active flow separation control.
The motivation for this study stems from the need for more robust and efficient flow control strategies capable of maintaining high lift coefficients near stall conditions, where traditional control methods tend to lose effectiveness. While several studies have applied MPC or RL separately, few have systematically compared their performance in a challenging aerodynamic control scenario. The novelty of this work lies in the development and comparison of adaptive MPC and RL-based techniques for real-time, closed-loop control of a NACA 4412 airfoil under separated flow conditions, using a dual-point plasma actuator configuration initially introduced by the authors in earlier studies. Notably, the RL methods employed here are trained entirely online and within the simulation environment, without relying on pre-generated datasets. This distinguishes our approach from conventional data-driven strategies, which typically depend on offline training using fixed databases. This study provides valuable insight into control stability, performance trade-offs, and the adaptability of modern learning-based control approaches in aerodynamic applications.
@@
The rest of the article is structured as follows: the Flow Control Actuator section introduces the concept of dual-point excitation using plasma actuators. The underlying computational framework, including the governing equations, plasma actuator model, computational grid, and numerical methods used, are explained in the Computational Procedure section. In the Flow Control Algorithms section, the various algorithms are investigated: adaptive MPC, TDRL, DQL, and signal processing integrated with DQL, each described with their respective algorithms, governing equations, and code implementations. The Results and Discussion section presents a thorough analysis of the findings, highlighting the comparative performance of the different control strategies and discussing their implications for future research and practical applications. Finally, the Conclusions section provides a summary of the study’s key findings.
@@ 2 Flow Control Actuator
Active flow separation control can generally be classified into two principal approaches: direct and indirect. In the direct method, actuators such as wall jets [33, 34] inject high-momentum fluid directly into the near-wall region. This process energizes the boundary layer, enabling fluid particles to resist adverse pressure gradients more effectively, thus delaying or suppressing flow separation. The indirect approach, on the other hand, utilizes unsteady excitation of the separated shear layer, typically achieved through pulsating actuators. The fundamental mechanism involves the generation and propagation of disturbances and vortices which enhance the entrainment of high-momentum fluid into the separated region. This increased mixing accelerates the replacement of low-energy fluid near the wall with more energetic fluid from the free stream, ultimately contributing to the postponement of separation [35, 36].
Over the years, numerous actuator technologies have been investigated for applying periodic excitation to separated shear layers. Among these are synthetic jets, combustion-powered actuators, microfiber composite piezoelectric actuators, and plasma actuators. Notably, dielectric barrier discharge (DBD) plasma actuators have attracted growing interest due to several advantageous characteristics. These include their fully electric operation, absence of moving components, suitability for integration into low-profile surfaces like airfoils, rapid response to unsteady inputs, low weight, and relatively low power consumption [37, 38]. Despite these benefits, DBD plasma actuators also exhibit certain limitations. These include susceptibility to electromagnetic interference, low efficiency in converting electrical energy to mechanical momentum, and limited induced flow velocity, typically less than 8 m/s [39, 40].
@@
Figure 1: (a) Schematic of a DBD plasma actuator and (b) Schematic of dual-point excitation technique (actuators are indicated with red marks) [42]. Source: Authors’ own work.
In previous work by the authors [43], the control of separated shear layers using DBD plasma actuators on a NACA 0015 airfoil was explored. The findings highlighted the actuator’s success in stall prevention and its positive impact on lift-to-drag ratios. Li et al. [44] employed plasma co-flow jets on a NACA 0025 airfoil, which significantly delayed flow separation through enhanced mixing shear layers. Samimy et al. [45] extended active flow control applications to high-speed flows using advancements in plasma actuators to manipulate Kelvin-Helmholtz instabilities. Additionally, Gross and Fasel [46] studied separation control on a NACA 643-618 airfoil using plasma pulsed vortex generator jets, demonstrating high efficiency through linear stability analysis.
Controlling flow separation through plasma actuators, particularly by exciting the separated shear layer, is highly sensitive to the choice of excitation frequency. Research indicates that a free shear layer is most responsive to perturbations near its natural vortex shedding frequency, leading many studies to select frequencies within this range for optimal separation control [47, 48, 49]. However, some research favors higher excitation frequencies [25, 50]. Another critical factor is the location of actuation, typically applied near the leading edge or the suction-side separation point to influence the upper separated shear layer [42]. Ebrahimi and Hajipour [36] introduced the concept of dual-point excitation, as shown in Fig. 1b, where both the suction side and trailing-edge shear layers are simultaneously excited using DBD plasma actuators. By exploiting the interaction between the two shear layers, their method reduces the separation region and attenuates unsteady aerodynamic load fluctuations under stall conditions, demonstrating superior control performance compared to conventional single-point excitation methods. In this research, we apply the same dual-point excitation method using DBD plasma actuators as the foundation of the flow control strategy, while the appropriate excitation frequency is treated as a parameter to be determined by the control algorithms.
@@ 3.1 Governing equations
The numerical flow simulations are performed by analyzing a two-dimensional incompressible viscous Newtonian fluid with constant properties. The governing equations employed for the Scale-Adaptive Simulation (SAS) turbulence model are based on the Reynolds-Averaged Navier-Stokes (RANS) equations, modified to include a scale-adaptive term that allows the model to transition between RANS and Large Eddy Simulation (LES) behavior based on the local flow conditions [51]. The continuity and momentum equations in Cartesian coordinates are:
@@ 3.2 Plasma actuator model
$E_{0}$ represents the maximum electric field strength within the triangular region, and the constants $k_{1}$ and $k_{2}$ are determined to ensure that the field strength at the boundary between the plasma and air (along the hypotenuse of the triangle) equals the breakdown value. Consequently, the body force within the plasma zone is given by:
@@
In this research, the parameters of the plasma actuator are consistent with those used by Shyy et al. [52]. The exposed and enclosed electrode lengths are 0.5 mm and 3 mm, respectively. The height of the electrodes is 0.1 mm, and the gap between them in the streamwise direction is 0.25 mm, as illustrated schematically in Fig. 2. The applied voltage is 9 kV (root mean square), with a frequency of 3 kHz. Accordingly, the dimensions of the plasma actuation zone are specified as $a=0.0015\,\text{m}$ in height and $b=0.003\,\text{m}$ in length. The numerical model of the DBD plasma actuator has been validated by comparing it with the work of Shy et al. [52]. They assumed that plasma actuators operate in a steady, continuous mode; thus, the body force is modulated based on the duty cycle and excitation frequency to simulate unsteady actions. In this research, a duty cycle of 50% and excitation frequencies ranging from $f=0$ to 400 Hz, corresponding to a normalized frequency of $F^{+}=fc/U_{\infty}$ in the range of 0 to 11, are employed.
Figure 2: Triangular zone associated with plasma formation [36]. Source: Authors’ own work.
@@ 3.3 Computational grid
The computational test case is based on a NACA 4412 airfoil with a chord length $c=0.4\,\text{m}$, $15^{\circ}$ angle of attack, and a Reynolds number $Re=4.0\times 10^{5}$. As depicted in Fig. 3, the computational domain extends $20c$ downstream, $12.5c$ upstream and $12.5c$ in the direction normal to the chord. The airfoil surface is modeled as an adiabatic wall with a no-slip boundary condition. A freestream with a constant velocity of 14.6 m/s enters the domain from the upstream and bottom boundaries, while the downstream and top boundaries are defined as pressure outlets with ambient pressure.
Figure 3: Side view of the computational domain. Source: Authors’ own work.
@@
Following the previous work of authors [36], the novel concept of dual excitation technique is utilized to control the separated shear layers. Two plasma zones are implemented in the grid: one on the suction side of the airfoil just downstream of the separated shear layer, and the other on the pressure side at the airfoil’s trailing edge. The dimensions of these zones are selected based on the work by Shy et al. [52], with the plasma body force applied only within these regions. The suction side actuator is located at $x=0.35c$, while the pressure side actuator is positioned at $x=0.99c$. The grid’s streamwise resolution in the plasma zones is chosen carefully to satisfy the accuracy requirements for predicting plasma wall jets and flow separation. Fig. 4 illustrates a detailed view of the plasma zones encircled in blue and the computational grid around the airfoil.
Figure 4: Detailed view of computational grid around airfoil and the plasma zones. Source: Authors’ own work.